# Efficiently learning depth-3 circuits via quantum agnostic boosting

### v1

@@ 3.2 Algorithm
Input: $\varepsilon\in(0,1)$, copies of $|\psi\rangle$, weak learner ${\mathcal{A}}_{\mathsf{WAL}}$ (Def. 3.1) with promise $\eta(\tau)=\eta_{1}\tau^{\eta_{2}}$. Output: List of parity states $L=\{|\chi_{i}\rangle\}_{i\in[\kappa]}$, coefficients $B=\{\beta_{i}\}_{i\in[\kappa]}$ 1 22mm Set error parameters $\varepsilon_{s}=(2/3)^{1/\eta_{2}+1}\varepsilon^{2}/16$ and $\varepsilon_{p}=\varepsilon/2$. /* Stage 1: Structure learning (Theorem 3.7) */ 3 Set $|\psi_{1}\rangle=|\psi\rangle$, $\alpha_{1}=1$, $L=\varnothing$. 4 Set parameter $\eta=\eta(\varepsilon_{s})$ with $\eta(\cdot)$ being the promise of ${\mathcal{A}}_{\mathsf{WAL}}$ (Theorem 3.2). 5 Set $t_{\max}=1/\eta(\varepsilon_{s})^{2}$, $\delta^{\prime}=\delta/(3t_{\mathrm{max}})$, $\kappa=0$. 6[2mm] for $t=1$ to $t_{\max}$ do 7  8  2mm Run the weak agnostic learner ${\mathcal{A}}_{\mathsf{WAL}}$ on $S_{\mathsf{WAL}}$ copies of $|\psi_{t}\rangle$ to learn a parity state $|\chi_{t}\rangle$. 9   Run SWAP test on $O(1/\eta^{2}\log(t_{\mathrm{max}}/\delta))$ copies of $|\psi_{t}\rangle,|\chi_{t}\rangle$ such that with probability $\geq 1-\delta^{\prime}$, one obtains an $\eta/2$ approximation of $|\langle\psi_{t}|\chi_{t}\rangle|^{2}$. Call the estimate $\nu_{t}$. 10 if $\nu_{t}<\eta$ then break loop. 11   Update $L\leftarrow L\cup\{|\chi_{t}\rangle\}$ and $\kappa\leftarrow\kappa+1$. 12   Set $_{T(t)}=\sumop\displaylimits_{i=1}^{t}|\chi_{i}\rangle\langle\chi_{i}|$. 13   Let $\widehat{\alpha}_{t+1}^{2}$ be an $\varepsilon_{s}/2$ approximation of $\alpha_{t+1}^{2}:=\mbox{$\parallel{(\mathbb{I}-{}_{T(t)})|\psi\rangle}\parallel$}_{2}^{2}$ by measuring $|\psi\rangle$ in the basis $\{\mathbb{I}-{}_{T(t)},{}_{T(t)}\}$, $O(1/\varepsilon_{s}^{2}\log(1/\delta^{\prime}))$ many times. 14 if $\widehat{\alpha}_{t+1}^{2}<\varepsilon_{s}$ then break loop. 15   Prepare $S_{\mathsf{WAL}}$ copies of $|\psi_{t+1}\rangle=(\mathbb{I}-{}_{T(t)})|\psi\rangle/\alpha_{t+1}$ by measuring $O(S_{\mathsf{WAL}}/\varepsilon_{s}\log(1/\delta^{\prime}))$ copies of $|\psi\rangle$ in the basis $\{\mathbb{I}-{}_{T(t)},{}_{T(t)}\}$ and post-selecting for the first outcome. 16  17 2mm /* Stage 2: Parameter learning (Theorem 3.12) */ 18 Set error parameters $\upsilon_{1}=(\varepsilon_{p}\cdot\eta)/(63\kappa)$, $\upsilon_{2}=(\varepsilon_{p}\cdot\sqrt{\eta})/(18\kappa)$, $\upsilon^{\prime}=(\varepsilon_{p}\cdot\sqrt{\eta})/(36\kappa)$. 19 Estimate $\xi_{1}$ of $|\langle\psi|\chi_{1}\rangle|$ using the SWAP test up to error $\upsilon_{1}$. 20 Estimate $\xi_{j}$ of $|\langle\psi|\chi_{j}\rangle|$ for all $j\geq 2$ using the SWAP test up to error $\upsilon_{2}$. 21 for $j=2$ to $\kappa$ do 22  23  2mm Prepare copies of $|\chi_{j}^{R}\rangle=(|\chi_{1}\rangle+|\chi_{j}\rangle)/\sqrt{2}$ and $|\chi_{j}^{I}\rangle=(|\chi_{1}\rangle+i|\chi_{j}\rangle)/\sqrt{2}$ which are promised to be stabilizer states, using Lemma 3.6. 24   Estimate $\gamma_{j}^{R}$ of $|\langle\chi_{j}^{R}|\psi\rangle|$ using SWAP test up to error $\upsilon^{\prime}$. 25   Estimate $\gamma_{j}^{I}$ of $|\langle\chi_{j}^{I}|\psi\rangle|$ using SWAP test up to error $\upsilon^{\prime}$. 26   Set $a_{j}=\frac{2(\gamma_{j}^{R})^{2}-\xi_{1}^{2}-\xi_{j}^{2}}{2\xi_{1}},\kern 5.0pt\text{ and }\kern 5.0ptb_{j}=\frac{2(\gamma_{j}^{I})^{2}-\xi_{1}^{2}-\xi_{j}^{2}}{2\xi_{1}}.$ 27   Set $\widehat{\beta}_{j}=a_{j}+ib_{j}$. 28  29Set $\widehat{\beta}_{j}\leftarrow\widehat{\beta}_{j}/\beta$, where $\beta=\mbox{$\parallel{\widehat{\beta}}\parallel$}_{2}$. return List of $t_{max}$ parities $L=\{|\chi_{i}\rangle\}_{i}$ and their coefficients $B=\{\widehat{\beta}_{i}\}_{i}$. Algorithm 1 Quantum agnostic boosting
@@ 4 Learning algorithms
In this section, we show how the protocol of quantum agnostic boosting (Algorithm 1) can be utilized for improper agnostic learning of decision trees, juntas and DNFs. Finally, we will give a learning protocol of depth-$3$ circuits, based on boosting, in the PAC learning model under the uniform distribution given access to quantum examples.
@@ 4.3 Agnostic learning DNFs
where $|\widehat{\phi}\rangle=\sumop\displaylimits_{i=1}^{\kappa}\beta_{i}|\phi_{i}\rangle$ with $\beta\in{\mathcal{B}}_{\infty}^{k}$ are coefficients for the parities $|\phi_{i}\rangle$ and $\kappa=\operatorname{poly}((s/\varepsilon)^{\log\log s/\varepsilon})$.
@@
$\sumop\displaylimits_{T\in\mathcal{T}}\widehat{f}(T)^{2}\geq 1-\gamma.$
@@
Then, there exists an algorithm that with probability $\geq 1-\delta$, outputs a parity state $|\phi\rangle\in{\cal S}_{{\mathcal{C}}_{\mathsf{Par}}}$ satisfying
$|\langle\psi|\phi\rangle|^{2}\geq\tau/s^{O(\log\log s)}-\varepsilon.$
@@
$\widetilde{O}\Big((s/\tau)^{O(\log\log s/\tau)}/(\tau\cdot\varepsilon^{2})\cdot\log(1/\delta))\Big)$
@@
$\widetilde{O}\Big(n\cdot(s/\tau)^{O(\log\log s/\tau)}/(\tau\cdot\varepsilon^{2})\cdot\log(1/\delta)\Big).$
@@
${\cal F}_{\mathsf{Par}}(|\psi\rangle)\geq(\sqrt{\tau}-\sqrt{\gamma})^{2}\cdot(s/\gamma)^{-O(\log\log s/\gamma)}\geq\tau\cdot(s/\tau)^{-O(\log\log s/\tau)},$
where we let $\gamma=\tau/8$. So by applying our agnostic parity learning algorithm (Lemma 4.2) with the lower bound there set to $\tau\cdot(s/\tau)^{-O(\log\log s/\tau)}$, we can learn a state $|\phi\rangle\in{\cal S}_{{\mathcal{C}}_{\mathsf{Par}}}$ such that
$|\langle\psi|\phi\rangle|^{2}\geq\tau\cdot(s/\tau)^{-O(\log\log s/\tau)}-\varepsilon.$
with probability $\geq 1-\delta$. The corresponding sample and complexity time complexity follows from Theorem 4.1 for $\varepsilon<\tau/s^{O(\log\log(s))}$, completing the proof. ∎
@@
We instantiate the agnostic boosting algorithm (Algorithm 1), as in Theorem 3.2, using the weak agnostic learner ${\mathcal{A}}_{\mathsf{WAL}}$ for ${\mathcal{C}}_{\textsf{DNF}(s)}$ from Lemma 4.9. The corresponding promise is $\eta(\tau)=\tau/(2s^{\star})$ with $s^{\star}=(s/\tau)^{\log\log s/\tau}$, $\eta_{1}=1/(2s^{\star})$ and $\eta_{2}=1$ (as defined in Theorem 3.2 for $\varepsilon$ of Lemma 4.2 set to $\tau/(2s^{\star})$). The corresponding sample complexity is $S_{\mathsf{WAL}}=\widetilde{O}((s/\varepsilon_{s})^{\log\log(s/\varepsilon_{s})}/\varepsilon_{s}^{3}\cdot\log 1/\delta)$ and $T_{\mathsf{WAL}}=\widetilde{O}(n(s/\varepsilon_{s})^{\log\log(s/\varepsilon_{s})}/\varepsilon_{s}^{3}\cdot\log 1/\delta)$ for error instantiated as $\varepsilon_{s}$ in Algorithm 1 (Theorem 3.7). The output of Theorem 3.2 is then a strong (improper) agnostic learner consuming $\operatorname{poly}((s/\varepsilon)^{\log\log(s/\varepsilon)},1/\varepsilon,\log(1/\delta))$ copies and $\operatorname{poly}(n,(s/\varepsilon)^{\log\log(s/\varepsilon)},1/\varepsilon,\log(1/\delta))$ time. ∎
@@ 4.4 PAC learning depth-3 circuits
In this section, we finally show how to quantum PAC learn depth-$3$ circuits. As mentioned earlier in the introduction (Section 1), the current state-of-the-art algorithm for learning depth-$3$ circuits in the PAC model with only classical examples has a time complexity of $n^{O(\log n)}$. Here, we show how to quantum PAC learn these circuits in $n^{O(\log\log n)}$ time. In order to prove our result, we will use the well-known discriminator lemma by Hajnal et al. [HMP+93] (which we reprove below specialized to our setting). Using our discriminator at each step of our boosting algorithm (in order to construct our weak learner), we are able to show that depth-$3$ circuits are learnable in the quantum PAC model with the desired time complexity as stated above.
@@ 4.4.2 Learning algorithm
$\operatorname{poly}(n,t,(s/\varepsilon)^{\log\log s/\varepsilon},\log(1/\delta)).$
@@
Consider the context of Theorem 4.14. Let ${\mathcal{C}}_{\textsf{DNF}(s)}$ be the class of interest and ${\mathcal{A}}_{\mathsf{WAL}}$ of Lemma 4.9 be the weak agnostic learner. Suppose we apply agnostic boosting (Algorithm 1 and Theorem 3.7) to $|\psi_{f}\rangle$ then the following is true. If $|\alpha_{t}|^{2}\geq\varepsilon$, then ${\cal F}_{{\mathcal{C}}}(|\psi_{t}\rangle)\geq\varepsilon/(4t^{2})$.
@@
Suppose as part of structure learning of the agnostic boosting algorithm (Algorithm 1), we have carried out t1t-1 iterations so far and learned the parity states $\{|\phi_{i}\rangle\}_{i\in[t-1]}$. Then, the residual quantum state $|\psi_{t}\rangle$ in the $t$th iteration can be expressed as
$|\psi_{t}\rangle=\frac{1}{\alpha_{t}}\Big(|\psi_{f}\rangle-{}_{T(t-1)}|\psi_{f}\rangle\Big)=\frac{1}{\alpha_{t}}\Big(|\psi_{f}\rangle-\sumop\displaylimits_{i=1}^{t-1}\beta_{i}|\phi_{i}\rangle\Big),$
where $T(t-1)=\mathrm{span}(\{|\phi_{i}\rangle\}_{i\in[t-1]})$, $\beta_{i}=\langle\phi_{i}|\psi_{f}\rangle$ (Eq. (8)) and $\alpha_{t}=\sqrt{1-\sumop\displaylimits_{i=1}^{t}\beta_{i}^{2}}$ (Fact 3.3). Let us define the distribution
$D_{f,t}(x)=\Big|f(x)-\sumop\displaylimits_{i=1}^{t}\beta_{i}\chi_{i}(x)\Big|\cdot 2^{-n}\cdot{}^{-1}$,
@@
$=2^{-n}\cdot{}^{-1}\cdot\sumop\displaylimits_{x}\left|f(x)-\sumop\displaylimits_{i=1}^{t}\beta_{i}\chi_{i}(x)\right|\cdot f(x)\cdot c(x)$
$=2^{-n}\cdot{}^{-1}\cdot\sumop\displaylimits_{x}\left(f(x)-\sumop\displaylimits_{i=1}^{t}\beta_{i}\chi_{i}(x)\right)\cdot c(x)$
@@
where the third equality used the fact that, for $a\in\{-1,1\}$ and $b\in[-1,1]$, we have that $|a-b|=a\cdot(a-b)$. [^16] To see this note that if $a=1$ then $a\geq b$ and $|a-b|=a-b=1-b=a\cdot(a-b).$ If a=1,a=-1, then $a\leq b$ and $|a-b|=b-a=b+1=a\cdot(a-b).$ One can then again use Theorem 4.12 to show that
@@
${\cal F}_{{\mathcal{C}}}(|\psi_{t}\rangle)\geq|\langle\psi_{t}|\psi_{c}\rangle|^{2}\geq\varepsilon/(4t^{2}).$
@@
Let $f$ be as in the theorem statement. We will employ the quantum agnostic boosting algorithm (Algorithm 1) on input of copies of $|\psi_{f}\rangle$ against the class ${\cal C}_{\textsf{DNF}(s)}$ and use the weak agnostic learner ${\mathcal{A}}_{\mathsf{WAL}}$ of Lemma 4.9.171717Our notation here is similar to the one in Section 3. Suppose we are at the end of the $t$th iteration of the agnostic boosting algorithm. We first observe that by Claim 4.15, if $\alpha_{t+1}^{2}\geq\varepsilon$, then ${\cal F}_{{\mathcal{C}}_{\textsf{DNF}(s)}}(|\psi_{t+1}\rangle)\geq\varepsilon/(4t^{2})$. If we were to then use ${\mathcal{A}}_{\mathsf{WAL}}$ of Lemma 4.9, we are guaranteed to learn a parity state $|\phi_{t+1}\rangle$ that satisfies $|\langle\chi_{t+1}|\psi_{t+1}\rangle|^{2}\geq\varepsilon/(4t^{2}s^{\star})$ where we have denoted $s^{\star}=(s/\varepsilon)^{\log\log s/\varepsilon}$. Thus, we no longer need to check the fidelity of the residual state with ${\cal C}_{\textsf{DNF}(s)}$. We now present the simplified stage $1$ of the algorithm in Algorithm 3
Input: Parameters $s,t$, $\varepsilon\in(0,1)$, copies of $|\psi\rangle$, weak agnostic learner ${\mathcal{A}}_{\mathsf{WAL}}$ of Lemma 4.9 Output: List of parities $L=\{|\chi_{i}\rangle\}_{i\in[\kappa]}$ 1 2Set error parameter $\varepsilon_{s}=\varepsilon/9$. 3 Set $|\psi_{1}\rangle=|\psi\rangle$, $\alpha_{1}=1$, $L=\varnothing$. 4 Set parameter $\eta=\eta(\varepsilon_{s})$ with $\eta(\cdot)$ being the promise of Lemma 4.9. (Theorem 3.2). 5 Set $t_{\max}=1/\eta(\varepsilon_{s})^{2}$, $\delta^{\prime}=\delta/(3t_{\mathrm{max}})$, $\kappa=0$. 6[2mm] for $t=1$ to $t_{\max}$ do 7  8  2mm Run the weak agnostic learner ${\mathcal{A}}_{\mathsf{WAL}}$ on $S_{\mathsf{WAL}}$ copies of $|\psi_{t}\rangle$ to learn a parity state $|\chi_{t}\rangle$. 9   Update $L\leftarrow L\cup\{|\chi_{t}\rangle\}$ and $\kappa\leftarrow\kappa+1$. 10   Set $_{T(t)}=\sumop\displaylimits_{i=1}^{t}|\chi_{i}\rangle\langle\chi_{i}|$. 11   Let $\widehat{\alpha}_{t+1}^{2}$ be an $\varepsilon_{s}/2$ approximation of $\alpha_{t+1}^{2}:=\mbox{$\parallel{(\mathbb{I}-{}_{T(t)})|\psi\rangle}\parallel$}_{2}^{2}$ by measuring $|\psi\rangle$ in the basis $\{\mathbb{I}-{}_{T(t)},{}_{T(t)}\}$, $O(1/\varepsilon_{s}^{2}\log(1/\delta^{\prime}))$ many times. 12 if $\widehat{\alpha}_{t+1}^{2}<\varepsilon_{s}$ then break loop. 13   Prepare $S_{\mathsf{WAL}}$ copies of $|\psi_{t+1}\rangle=(\mathbb{I}-{}_{T(t)})|\psi\rangle/\alpha_{t+1}$ by measuring $O(S_{\mathsf{WAL}}/\varepsilon_{s}\log(1/\delta^{\prime}))$ copies of $|\psi\rangle$ in the basis $\{\mathbb{I}-{}_{T(t)},{}_{T(t)}\}$ and post-selecting for the first outcome. 14  15 2mm return List of $\kappa$ parity states $L=\{|\chi_{i}\rangle\}_{i}$. Algorithm 3 Structure learning for depth-$3$ circuits

### v2

@@ 3.2 Algorithm
Input: $\varepsilon\in(0,1)$, copies of $|\psi\rangle$, weak learner ${\mathcal{A}}_{\mathsf{WAL}}$ (Def. 3.1) with promise $\eta(\tau)=\eta_{1}\tau^{\eta_{2}}$. Output: List of parity states $L=\{|\chi_{i}\rangle\}_{i\in[\kappa]}$, coefficients $B=\{\beta_{i}\}_{i\in[\kappa]}$ 1 22mm Set error parameters $\varepsilon_{s}=(2/3)^{1/\eta_{2}+1}\varepsilon^{2}/16$ and $\varepsilon_{p}=\varepsilon/2$. /* Stage 1: Structure learning (Theorem 3.7) */ 3 Set $|\psi_{1}\rangle=|\psi\rangle$, $\alpha_{1}=1$, $L=\varnothing$. 4 Set parameter $\eta=\eta(\varepsilon_{s})$ with $\eta(\cdot)$ being the promise of ${\mathcal{A}}_{\mathsf{WAL}}$ (Theorem 3.2). 5 Set $t_{\max}=4/(\varepsilon_{s}\eta(\varepsilon_{s}))$, $\delta^{\prime}=\delta/(3t_{\mathrm{max}})$, $\kappa=0$. 6[2mm] for $t=1$ to $t_{\max}$ do 7  8  2mm Run the weak agnostic learner ${\mathcal{A}}_{\mathsf{WAL}}$ on $S_{\mathsf{WAL}}$ copies of $|\psi_{t}\rangle$ to learn a parity state $|\chi_{t}\rangle$. 9   Run SWAP test on $O(1/\eta^{2}\log(t_{\mathrm{max}}/\delta))$ copies of $|\psi_{t}\rangle,|\chi_{t}\rangle$ such that with probability $\geq 1-\delta^{\prime}$, one obtains an $\eta/2$ approximation of $|\langle\psi_{t}|\chi_{t}\rangle|^{2}$. Call the estimate $\nu_{t}$. 10 if $\nu_{t}<\eta$ then break loop. 11   Update $L\leftarrow L\cup\{|\chi_{t}\rangle\}$ and $\kappa\leftarrow\kappa+1$. 12   Set $_{T(t)}=\sumop\displaylimits_{i=1}^{t}|\chi_{i}\rangle\langle\chi_{i}|$. 13   Let $\widehat{\alpha}_{t+1}^{2}$ be an $\varepsilon_{s}/2$ approximation of $\alpha_{t+1}^{2}:=\mbox{$\parallel{(\mathbb{I}-{}_{T(t)})|\psi\rangle}\parallel$}_{2}^{2}$ by measuring $|\psi\rangle$ in the basis $\{\mathbb{I}-{}_{T(t)},{}_{T(t)}\}$, $O(1/\varepsilon_{s}^{2}\log(1/\delta^{\prime}))$ many times. 14 if $\widehat{\alpha}_{t+1}^{2}<\varepsilon_{s}$ then break loop. 15   Prepare $S_{\mathsf{WAL}}$ copies of $|\psi_{t+1}\rangle=(\mathbb{I}-{}_{T(t)})|\psi\rangle/\alpha_{t+1}$ by measuring $O(S_{\mathsf{WAL}}/\varepsilon_{s}\log(1/\delta^{\prime}))$ copies of $|\psi\rangle$ in the basis $\{\mathbb{I}-{}_{T(t)},{}_{T(t)}\}$ and post-selecting for the first outcome. 16  17 2mm /* Stage 2: Parameter learning (Theorem 3.12) */ 18 Set error parameters $\upsilon_{1}=(\varepsilon_{p}\cdot\eta)/(63\kappa)$, $\upsilon_{2}=(\varepsilon_{p}\cdot\sqrt{\eta})/(18\kappa)$, $\upsilon^{\prime}=(\varepsilon_{p}\cdot\sqrt{\eta})/(36\kappa)$. 19 Estimate $\xi_{1}$ of $|\langle\psi|\chi_{1}\rangle|$ using the SWAP test up to error $\upsilon_{1}$. 20 Estimate $\xi_{j}$ of $|\langle\psi|\chi_{j}\rangle|$ for all $j\geq 2$ using the SWAP test up to error $\upsilon_{2}$. 21 for $j=2$ to $\kappa$ do 22  23  2mm Prepare copies of $|\chi_{j}^{R}\rangle=(|\chi_{1}\rangle+|\chi_{j}\rangle)/\sqrt{2}$ and $|\chi_{j}^{I}\rangle=(|\chi_{1}\rangle+i|\chi_{j}\rangle)/\sqrt{2}$ which are promised to be stabilizer states, using Lemma 3.6. 24   Estimate $\gamma_{j}^{R}$ of $|\langle\chi_{j}^{R}|\psi\rangle|$ using SWAP test up to error $\upsilon^{\prime}$. 25   Estimate $\gamma_{j}^{I}$ of $|\langle\chi_{j}^{I}|\psi\rangle|$ using SWAP test up to error $\upsilon^{\prime}$. 26   Set $a_{j}=\frac{2(\gamma_{j}^{R})^{2}-\xi_{1}^{2}-\xi_{j}^{2}}{2\xi_{1}},\kern 5.0pt\text{ and }\kern 5.0ptb_{j}=\frac{2(\gamma_{j}^{I})^{2}-\xi_{1}^{2}-\xi_{j}^{2}}{2\xi_{1}}.$ 27   Set $\widehat{\beta}_{j}=a_{j}+ib_{j}$. 28  29Set $\widehat{\beta}_{j}\leftarrow\widehat{\beta}_{j}/\beta$, where $\beta=\mbox{$\parallel{\widehat{\beta}}\parallel$}_{2}$. return List of $\kappa$ parity states $L=\{|\chi_{i}\rangle\}_{i}$ and their coefficients $B=\{\widehat{\beta}_{i}\}_{i}$. Algorithm 1 Quantum agnostic boosting
@@ 4 Learning algorithms
In this section, we show how quantum agnostic boosting (Algorithm 1) can be utilized for improper agnostic learning of decision trees, juntas and DNFs. Finally, we will give a learning protocol of depth-$3$ circuits, based on boosting, in the uniform PAC model given quantum examples.
@@ 4.3 Agnostic learning DNFs
where $|\widehat{\phi}\rangle=\sumop\displaylimits_{i=1}^{\kappa}\beta_{i}|\phi_{i}\rangle$ with coefficients $\beta\in{\mathcal{B}}_{\infty}^{k}$ for the parities $|\phi_{i}\rangle$ and $\kappa=\operatorname{poly}((s/\varepsilon)^{\log\log(s/\varepsilon)\cdot\log(1/\varepsilon)})$.
@@
$\sumop\displaylimits_{T\in\mathcal{T}_{\gamma}}\widehat{f}(T)^{2}\geq 1-\gamma.$
@@
Let $s^{*}=(s/\tau)^{O(\log\log(s/\tau)\cdot\log(1/\tau))}$. Then, there exists an algorithm that with probability $\geq 1-\delta$, outputs a parity state $|\phi\rangle\in{\cal S}_{{\mathcal{C}}_{\mathsf{Par}}}$ satisfying
$|\langle\psi|\phi\rangle|^{2}\geq\tau/s^{*}-\varepsilon.$
@@
$\widetilde{O}\Big(s^{*}/(\tau\cdot\varepsilon^{2})\cdot\log(1/\delta))\Big)$
@@
$\widetilde{O}\Big(n\cdot s^{*}/(\tau\cdot\varepsilon^{2})\cdot\log(1/\delta)\Big).$
@@
${\cal F}_{\mathsf{Par}}(|\psi\rangle)\geq(\sqrt{\tau}-\sqrt{\gamma})^{2}\cdot(s/\gamma)^{-O(\log\log(s/\gamma)\log(1/\gamma)}\geq\tau\cdot\underbrace{(s/\tau)^{-O(\log\log(s/\tau)\cdot\log(1/\tau))}}_{:=1/s^{*}},$
where we let $\gamma=\tau/8$. So by applying our agnostic parity learning algorithm (Lemma 4.2) with the lower bound there set to $\tau/s^{*}$, we can learn a state $|\phi\rangle\in{\cal S}_{{\mathcal{C}}_{\mathsf{Par}}}$ such that
$|\langle\psi|\phi\rangle|^{2}\geq\tau/s^{*}-\varepsilon.$
with probability $\geq 1-\delta$. The corresponding sample and complexity time complexity follows from Theorem 4.1 for $\varepsilon<\tau/s^{*}$, completing the proof. ∎
@@
We instantiate the agnostic boosting algorithm (Algorithm 1), as in Theorem 3.2, using the weak agnostic learner ${\mathcal{A}}_{\mathsf{WAL}}$ for ${\mathcal{C}}_{\textsf{DNF}(s)}$ from Lemma 4.9. The corresponding promise is $\eta(\tau)=\tau/(2s^{\star})$ with $s^{*}=(s/\tau)^{O(\log\log(s/\tau)\cdot\log(1/\tau))}$, $\eta_{1}=1/(2s^{\star})$ and $\eta_{2}=1$ (as defined in Theorem 3.2 for $\varepsilon$ of Lemma 4.2 set to $\tau/(2s^{\star})$). Note that the function $\tau/s^{*}$ is an increasing function of $\tau$ (as required by the boosting algorithm). The corresponding sample complexity is $S_{\mathsf{WAL}}=\widetilde{O}((s/\varepsilon_{s})^{\log\log(s/\varepsilon_{s})\log(1/\varepsilon_{s})}/\varepsilon_{s}^{3}\cdot\log 1/\delta)$ and $T_{\mathsf{WAL}}=\widetilde{O}(n(s/\varepsilon_{s})^{\log\log(s/\varepsilon_{s})\log(1/\varepsilon_{s})}/\varepsilon_{s}^{3}\cdot\log 1/\delta)$ for error instantiated as $\varepsilon_{s}$ in Algorithm 1 (Theorem 3.7). The output of Theorem 3.2 is then a strong (improper) agnostic learner consuming $\operatorname{poly}(s^{*},1/\varepsilon,\log(1/\delta))$ copies and $\operatorname{poly}(n,s^{*},1/\varepsilon,\log(1/\delta))$ time with $s^{*}=(s/\varepsilon)^{O(\log\log(s/\varepsilon)\cdot\log(1/\varepsilon))}$. ∎
@@ 4.4 PAC learning depth-3 circuits
In this section, we finally show how to quantum PAC learn depth-$3$ circuits. As mentioned earlier in the introduction (Section 1), the current state-of-the-art algorithm for learning depth-$3$ circuits in the PAC model with only classical examples has a time complexity of $n^{O(\log^{2}n)}$. Here, we show how to quantum PAC learn these circuits in $n^{O(\log n)}$ time. In order to prove our result, we will use the well-known discriminator lemma by Hajnal et al. [HMP+93] (which we reprove below specialized to our setting). Using our discriminator at each step of our boosting algorithm (in order to construct our weak learner), we are able to show that depth-$3$ circuits are learnable in the quantum PAC model with the desired time complexity as stated above.
@@ 4.4.2 Learning algorithm
$\operatorname{poly}(n,m,(s/\varepsilon)^{(\log\log s/\varepsilon)\cdot\log 1/\varepsilon},\log(1/\delta)).$
@@
Consider the context of Theorem 4.14. Let ${\mathcal{C}}_{\textsf{DNF}(s)}$ be the class of interest and ${\mathcal{A}}_{\mathsf{WAL}}$ of Lemma 4.9 be the weak agnostic learner. Suppose we apply agnostic boosting (Algorithm 1 and Theorem 3.7) to $|\psi_{f}\rangle$ then the following is true. If $|\alpha_{t}|^{2}\geq\varepsilon$, then ${\cal F}_{{\mathcal{C}}_{\textsf{DNF}(s)}}(|\psi_{t}\rangle)\geq\varepsilon/4m^{2}$.
@@
Suppose as part of structure learning of the agnostic boosting algorithm (Algorithm 1), we have carried out t1t-1 iterations so far and learned the parity states $\{|\chi_{i}\rangle\}_{i\in[t-1]}$. Then, the residual quantum state $|\psi_{t}\rangle$ in the $t$th iteration can be expressed as
$|\psi_{t}\rangle=\frac{1}{\alpha_{t}}\Big(|\psi_{f}\rangle-{}_{T(t-1)}|\psi_{f}\rangle\Big)=\frac{1}{\alpha_{t}}\Big(|\psi_{f}\rangle-\sumop\displaylimits_{i=1}^{t-1}\beta_{i}|\chi_{i}\rangle\Big),$
where $T(t-1)=\mathrm{span}(\{|\phi_{i}\rangle\}_{i\in[t-1]})$, $\beta_{i}=\langle\chi_{i}|\psi_{f}\rangle$ (Eq. (8)) and $\alpha_{t}=\sqrt{1-\sumop\displaylimits_{i=1}^{t-1}|\beta_{i}|^{2}}$ (Fact 3.3). Let us denote the states
$|\psi_{f}\rangle=\frac{1}{\sqrt{2^{n}}}\sumop\displaylimits_{x\in\{0,1\}^{n}}f(x)|x\rangle,\kern 5.0pt|\chi_{i}\rangle=\frac{1}{\sqrt{2^{n}}}\sumop\displaylimits_{x\in\{0,1\}^{n}}\chi_{i}(x)|x\rangle,$
where $f:\{0,1\}^{n}\rightarrow\{-1,1\}$ and $\chi_{i}(x)=(-1)^{\alpha_{i}\cdot x}$ for some $\alpha_{i}\in\{0,1\}^{n},\,\,\forall i\in[t-1]$. Let us define the distribution
$D_{f,t}(x)=\Big|f(x)-\sumop\displaylimits_{i=1}^{t-1}\beta_{i}\chi_{i}(x)\Big|\cdot 2^{-n}\cdot{}^{-1}$,
@@
$=2^{-n}\cdot{}^{-1}\cdot\sumop\displaylimits_{x}\left|f(x)-\sumop\displaylimits_{i=1}^{t-1}\beta_{i}\chi_{i}(x)\right|\cdot f(x)\cdot c(x)$
$=2^{-n}\cdot{}^{-1}\cdot\sumop\displaylimits_{x}\left(f(x)-\sumop\displaylimits_{i=1}^{t-1}\beta_{i}\chi_{i}(x)\right)\cdot c(x)$
@@
where the third equality used the fact that, for $a\in\{-1,1\}$ and $b\in[-1,1]$, we have that $|a-b|=a\cdot(a-b)$. [^16] To see this note that if $a=1$ then $a\geq b$ and $|a-b|=a-b=1-b=a\cdot(a-b).$ If a=1,a=-1, then $a\leq b$ and $|a-b|=b-a=b+1=a\cdot(a-b).$ One can then use Lemma 4.12 to show that
@@
${\cal F}_{{\mathcal{C}}}(|\psi_{t}\rangle)\geq|\langle\psi_{t}|\psi_{c}\rangle|^{2}\geq\varepsilon/4m^{2}.$
@@
Let $f$ be as in the theorem statement. We will employ the quantum agnostic boosting algorithm (Algorithm 1) on input of copies of $|\psi_{f}\rangle$ against the class ${\cal C}_{\textsf{DNF}(s)}$ and use the weak agnostic learner ${\mathcal{A}}_{\mathsf{WAL}}$ of Lemma 4.9.171717Our notation here is similar to the one in Section 3. Suppose we are at the end of the $t$th iteration of the agnostic boosting algorithm. We first observe that by Claim 4.15, if $\alpha_{t+1}^{2}\geq\varepsilon$, then ${\cal F}_{{\mathcal{C}}_{\textsf{DNF}(s)}}(|\psi_{t+1}\rangle)\geq\varepsilon/(4m^{2})$. If we were to then use ${\mathcal{A}}_{\mathsf{WAL}}$ of Lemma 4.9, we are guaranteed to learn a parity state $|\phi_{t+1}\rangle$ that satisfies $|\langle\chi_{t+1}|\psi_{t+1}\rangle|^{2}\geq\varepsilon/(4m^{2}s^{\star})$ where we have denoted $s^{\star}=(s/\varepsilon)^{(\log\log s/\varepsilon)\cdot\log 1/\varepsilon}$. Note that the function $\varepsilon/s^{*}$ is an increasing function of $\varepsilon$ (as required by the boosting algorithm). Thus, we no longer need to check the fidelity of the residual state with ${\cal C}_{\textsf{DNF}(s)}$. We now present the simplified stage $1$ of the algorithm in Algorithm 3
[…]